# Equilibrium Cycle: A “Dynamic” Equilibrium

### v2

@@ Abstract
equilibrium cycle,
non-cooperative games,
discontinuous games,
equilibrium concept,
oscillatory game dynamics,
curb sets,
@@ 1 Introduction
Game theory, as a mathematical framework, helps us to understand the behaviour of rational agents in strategic interactions. A fundamental concept in game theory is the Nash equilibrium (NE), which is considered as the outcome of the game, or indeed, “the meaning of the game” (Milionis et al., 2022). Formally, a Nash equilibrium represents an action profile satisfying the property that no player has a unilateral incentive to deviate from it.
However, there are well known issues with applying the Nash equilibrium to understand game dynamics. By game dynamics, we mean settings where players change their actions over time, seeking to improve their own payoff, in response to the actions of other players. Such game dynamics, which includes the special case of best response dynamics, do not always converge to a Nash equilibrium (Demichelis et al., 2003; Benaïm et al., 2012). From a dynamical systems standpoint, this is not surprising. After all, NEs are simply stationary points (a.k.a., equilibria) corresponding to a broad class of game dynamics, and dynamical systems do not in general converge to a stationary point. Indeed, it is not unusual for game dynamics to result in limit cycles (i.e., the action profile approaches a deterministic periodic cycle asymptotically), or even more chaotic behavior in higher dimensions; see Benaïm and Hirsch (1999); Ficici et al. (2005); Benaïm et al. (2012); Papadimitriou and Piliouras (2019). In such situations, the Nash equilibrium as a solution concept does not meaningfully capture the ‘outcome’ of the (dynamic) strategic interaction between the players.
In this paper, we introduce a novel solution concept, which we call the equilibrium cycle (EC), that seeks to capture the outcome of oscillatory game dynamics. Specifically, we restrict attention to pure (i.e., non-randomised) player strategies. The EC seeks to capture the limit set associated with a broad class of such (oscillatory) game dynamics. Crucially, the definition of the EC does not require the existence of best responses, and is therefore also applicable to discontinuous games, which arise naturally in various contexts (Dasgupta and Maskin, 1986a, b; Reny, 1999). [^1] We first discovered the equilibrium cycle while analyzing a certain pricing game between competing ride-hailing platforms (Walunj et al., 2023). This papers seeks to formalize this equilibrium notion in general terms, provide examples of equilibrium cycles in economics applications, and make connections with other classical equilibrium notions.
@@
To summarize, we see that better response dynamics in the visibility game described in Example 1 do not converge; rather, the action profile oscillates indefinitely in the set $[0,0.5]^{2}$. Moreover, this set $[0,0.5]^{2}$ of action profiles satisfies the following properties:
@@
As we show in Section 2, the set $[0,0.5]^{2}$ is an equilibrium cycle corresponding to the game in Example 1. Indeed, the EC is characterized via the above three defining properties: stability, unrest, and minimiality. Intuitively, the first property ensures that under game dynamics, once the action profile enters an EC, it remains in the EC. The second property ensures that the action profile oscillates within the EC indefinitely. The third property ensures that the EC characterization is tight, i.e., ‘irrelevant’ action profiles are not included within the set.
@@ 2 Equilibrium Cycle: Definition and examples
Consider a strategic form game ${\mathcal{N}},({\mathcal{A}}_{i})_{i\in{\mathcal{N}}},({\mathcal{U}}_{i})_{i\in{\mathcal{N}}}\right>,$ where for each $i,$ ${\mathcal{A}}_{i}$ is a nonempty set in a metric space.
@@
    [Stability] For any player $i$ and opponent action profile ${\bf a}_{-i}\in{\mathcal{E}}_{-i},$ there exists an action $a_{i}\in{\mathcal{E}}_{i}$ such that
    ${\mathcal{U}}_{i}(a_{i},{\bf a}_{-i})>{\mathcal{U}}_{i}(\tilde{a}_{i},{\bf a}_{-i})\quad\text{for all actions }\tilde{a}_{i}\in{\mathcal{A}}_{i}\setminus{\mathcal{E}}_{i}.$
@@
    [Unrest] For any action profile ${\bf a}\in{\mathcal{E}},$ there exists a player $i\in{\mathcal{N}},$ and an alternate action $a_{i}^{\prime}\in{\mathcal{E}}_{i}$ such that ${\mathcal{U}}_{i}(a_{i}^{\prime},{\bf a}_{-i})>{\mathcal{U}}_{i}(a_{i},{\bf a}_{-i}),$ and further,
    ${\mathcal{U}}_{i}(a_{i}^{\prime},{\bf a}_{-i})>{\mathcal{U}}_{i}(\tilde{a}_{i},{\bf a}_{-i})\quad\text{for all actions }\tilde{a}_{i}\in{\mathcal{A}}_{i}\setminus{\mathcal{E}}_{i}.$
@@
Lemma 2.
For the visibility game discussed in Example 1, the set $[0,0.5]^{2}$ is an equilibrium cycle.
@@
If $a_{-i}<0.5,$ then any choice of $a_{i}\in(a_{-i},0.5),$ which implies ${\mathcal{U}}_{i}(a_{i},a_{-i})=1-a_{i},$ satisfies (2). On the other hand, if $a_{-i}=0.5,$ then $a_{i}=0,$ which implies ${\mathcal{U}}_{i}(a_{i},a_{-i})=0.5,$ satisfies (2). This proves that $[0,0.5]^{2}$ satisfies the first condition for an EC.
Next, we prove that $[0,0.5]^{2}$ satisfies the second condition for an EC. Consider any $(a_{1},a_{2})\in[0,0.5]^{2}.$
@@
This proves that $[0,0.5]^{2}$ satisfies the second condition for an EC.
To establish the third condition, suppose, for the purpose of obtaining a contradiction, that ${\mathcal{F}}\subsetneq{\mathcal{E}}$ is a non-empty closed Cartesian product set satisfying the first two conditions for an EC. Define the diagonal ${\mathcal{D}}:=\{(x,x):x\in[0,0.5]\}$. Note that we cannot have ${\mathcal{D}}\subset{\mathcal{F}},$ since this would imply (given that ${\mathcal{F}}$ is a Cartesian product) that ${\mathcal{F}}={\mathcal{E}}.$ It follows that ${\mathcal{F}}\cap{\mathcal{D}}\subsetneq{\mathcal{D}}.$ Consider the following cases.
@@ 2. $(ii)$
    $\{x\in[0,b)\colon(x,x)\in{\mathcal{F}}\}\neq\emptyset:$ In this case, define
    $c:=\sup\{x:(x,x)\in{\mathcal{F}}\cap[0,b)^{2}\}.$
    By the construction of $b$ and $c$, it follows that $c<b,$ $(c,c)$ lies in ${\mathcal{F}},$ and $(x,x)\notin{\mathcal{F}}$ for all $x\in(c,b).$ Now, for any player $i,$ fix opponent’s action $a_{-i}=c.$ Then for $a_{i}\in(c,b),$
    ${\mathcal{U}}_{i}(a_{i},c)=1-a_{i}>{\mathcal{U}}_{i}(\tilde{a}_{i},c)\ \forall\ \tilde{a}_{i}\in{\mathcal{F}}_{i}.$
    This contradicts that ${\mathcal{F}}$ satisfies the first condition for an EC.
In summary, we have shown (via a contradiction-based argument) that there does not exist a closed Cartesian product ${\mathcal{F}}\subsetneq{\mathcal{E}}$ that satisfies the first two conditions of an EC. We conclude that $[0,0.5]^{2}$ is an EC for the visibility game in Example 1.
Interestingly, it is known that the visibility game of Example 1 admits a mixed Nash equilibrium supported on $[0,1-\nicefrac{{1}}{{e}}]^{2},$ where $e$ denotes Euler’s number; see Lotker et al. (2008). It is instructive to note that the support of the mixed NE differs from the EC, i.e., $[0,0.5]^{2}.$
@@
It can be shown that $[0,\nicefrac{{(N-1)}}{{N}}]^{N}$ is an EC for this $N$ player extension. Observe that the action set corresponding to each player in this EC grows as $N$ increases.
@@
Consider the following variation of the two-player Bertrand duopoly (see Osborne (2009)) where firms additionally incur a fixed operational cost. Formally, the firms have a fixed marginal cost of production $c>0$, and a fixed operational cost $O_{c}>0$ for participating in the market. Each player/firm $i$ has two options: it can either operate (incurring the cost $O_{c}$) by setting price $p_{i}\geq 0$, or choose not to operate, via the ‘opt-out’ action $p_{i}=n_{o}<0$ (the opt-out action is represented by a negative number for mathematical convenience), incurring no costs. Thus, the action space of each player is modeled as $[0,\infty)\cup\{n_{o}\}.$
@@
${\mathcal{E}}:=\left([p_{m}^{*},p_{b}]\cup\{n_{o}\}\right)^{2}$
@@
It is instructive to note that in the classical Bertrand duopoly, $O_{c}=0,$ resulting in the unique Nash equilibrium $(c,c).$ It is the (reasonable) introduction of a positive operational cost (that is not so large that it prohibits firms from operating at all, as ensured by our assumption) that results in the EC characterized in Lemma 3. Finally, we note that the linear demand assumption is not essential to the emergence of the EC; an analogous EC can be established under any smooth decreasing demand function (so long as a positive utility is possible). The proof of Lemma 3 is in Appendix A.
The Bertrand duopoly posits an extreme ‘all or nothing’ bifurcation of market demand between the firms. The following example, which generalizes a pricing game between competing ride-hailing platforms in Walunj et al. (2023), considers a ‘less extreme’ payoff discontinuity.
@@
The game in Example 3 also admits an EC, as shown in the following lemma; this generalizes Theorem 4 in Walunj et al. (2023). The proof is once again omitted, given its similarity to the proof of Lemma 2.
Lemma 4.
Consider the game described in Example 3. Let $c:=\operatorname*{arg\,max}_{a_{i}\in{\mathcal{A}}_{i}}g(a_{i}),$ and $b:=f^{-1}(g(b)).$ Then ${\mathcal{E}}=[b,c]^{2}$ is an equilibrium cycle of this game.
@@ 3.1 Curb sets in best response games
    To prove that $C$ satisfies the first property of an EC, consider any player $i$ and ${\bf a}_{-i}\in C_{-i}.$ Using the definition of the curb set (see (5)), against ${\bf a}_{-i},$ there exists a best response $a_{i}\in C_{i}$ such that
    ${\mathcal{U}}_{i}(a_{i},{\bf a}_{-i})>{\mathcal{U}}_{i}(\tilde{a}_{i},{\bf a}_{-i})\text{ for all }\tilde{a}_{i}\notin C_{i}$
    (indeed, note that no other outside action $\tilde{a}_{i}\notin C_{i}$ is a best response). Thus, the first property of EC is satisfied.
@@
    Consider any strategy profile ${\bf a}\in C$. Since $C$ is a non-trivial curb set, ${\bf a}$ is not an NE, and thus there exists at least one player (say player $i$) which stands to obtain a strictly better utility via unilateral deviation. Suppose $a_{i}^{\prime}$ is a best response of this player $i;$ and we have ${\mathcal{U}}_{i}(a_{i}^{\prime},{\bf a}_{-i})>{\mathcal{U}}(a_{i},{\bf a}_{-i})$. Further, by the definition of curb set, $a_{i}^{\prime}\in C_{i}$ and as before,
    ${\mathcal{U}}_{i}(a_{i}^{\prime},{\bf a}_{-i})>{\mathcal{U}}_{i}(\tilde{a}_{i},{\bf a}_{-i})\text{ for all }\tilde{a}_{i}\notin C_{i}.$
    Thus, $C$ satisfies the second property of an EC.
@@
    We prove that $C$ satisfies the third property of an EC via a contradiction-based argument. Specifically, suppose that there exists ${\mathcal{E}}\subsetneq C$ satisfying the first two properties of an EC. Consider any player $i$. Against any ${\bf a}_{-i}\in{\mathcal{E}}_{-i}$, from the first property of EC, there exists $a_{i}^{\prime}\in{\mathcal{E}}_{i}$ such that
    ${\mathcal{U}}_{i}(a_{i}^{\prime},{\bf a}_{-i})>{\mathcal{U}}_{i}(\tilde{a}_{i},{\bf a}_{-i})\text{ for all }\tilde{a}_{i}\notin{\mathcal{E}}_{i}.$
    This implies that the best responses of player $i$ must be contained in ${\mathcal{E}}_{i},$ which in turn implies that the set ${\mathcal{E}}$ contains all its best responses. This means that ${\mathcal{E}}\subsetneq C$ is also a curb set, thus contradicting the assumption that $C$ is a minimal curb set.

### v3

@@ Abstract
equilibrium cycle , non-cooperative games , discontinuous games , equilibrium concept , oscillatory game dynamics , curb sets
journal: Games and Economic Behavior
[…]
@@ 1 Introduction
Game theory, as a mathematical framework, helps us to understand the behaviour of rational agents in strategic interactions. A fundamental concept initially proposed in game theory is the Nash equilibrium (NE), which was considered as the outcome of the game, or indeed, “the meaning of the game” (Nash Jr, 1950; Milionis et al., 2022). Formally, a Nash equilibrium represents an action profile satisfying the property that no player has a unilateral incentive to deviate from it.
However, there are well known issues with applying the Nash equilibrium to understand game dynamics. By game dynamics, we mean settings where players change their actions over time, seeking to improve their own payoff, in response to the actions of other players. Such game dynamics, which includes the special case of best response dynamics, do not always converge to a Nash equilibrium (Demichelis et al., 2003; Milionis et al., 2022; Benaïm et al., 2012). From a dynamical systems standpoint, this is not surprising. After all, NEs are simply stationary points (a.k.a., equilibria) corresponding to a broad class of game dynamics, and dynamical systems do not in general converge to a stationary point. Indeed, it is not unusual for game dynamics to result in limit cycles (i.e., the action profile approaches a deterministic periodic cycle asymptotically), or even more chaotic behavior in higher dimensions; see Benaïm and Hirsch (1999); Ficici et al. (2005); Benaïm et al. (2012); Papadimitriou and Piliouras (2019); Milionis et al. (2022). In such situations, the Nash equilibrium as a solution concept does not meaningfully capture the ‘outcome’ of the (dynamic) strategic interaction between the players.
In this paper, we introduce a novel solution concept, which we call the equilibrium cycle (EC), that seeks to capture the outcome of oscillatory game dynamics. Specifically, we restrict attention to pure (i.e., non-randomised) player strategies. The EC seeks to capture the limit set associated with a broad class of such (oscillatory) game dynamics. Crucially, the definition of the EC does not require the existence of best responses, and is therefore also applicable to discontinuous games, which arise naturally in various contexts (Dasgupta and Maskin, 1986a, b; Reny, 1999). [^1] We first discovered the equilibrium cycle while analyzing a certain pricing game between competing ride-hailing platforms (Walunj et al., 2025b). This papers seeks to formalize this equilibrium notion in general terms, provide examples of equilibrium cycles in economics applications, and make connections with other classical equilibrium notions.
@@
To summarize, we see that better response dynamics (among pure strategies) in the visibility game described in Example 1 do not converge; rather, the action profile oscillates indefinitely in the set $0,0.5]^{2}$. Moreover, this set $0,0.5]^{2}$ of action profiles satisfies the following properties:
@@
As we show in Section 2, the set $0,0.5]^{2}$ is an equilibrium cycle corresponding to the game in Example 1. Indeed, the EC is characterized via the above three defining properties: stability, unrest, and minimiality. Intuitively, the first property ensures that under game dynamics, once the action profile enters an EC, it remains in the EC. The second property ensures that the action profile oscillates within the EC indefinitely. The third property ensures that the EC characterization is tight, i.e., ‘irrelevant’ action profiles are not included within the set.
Before we proceed with the formal definition of EC and subsequent analysis, we conclude this section with a discussion on the related literature.
1.1 Related Literature
The initial strands of literature analyzing strategic interactions among multiple agents, predominately describe the outcomes via static concepts, mainly the Nash Equilibrium (Nash Jr, 1950). After the initial quest, substantial work has examined dynamic strategic interactions that more closely reflect the real-world applications.
There are strands of literature that capture the scenarios where finitely many agents continually learn the strategies of other agents, in a bid to choose better strategies that provide them ‘optimal’ utilities. For example, when the agents primarily consider the most recent moves of the opponents, we have the well known best-response dynamics and when they track significant history of the same we have fictitious play (e.g., Brown (1951); Shapley (1963); Binmore et al. (2001)). Another branch of literature deals with large population games, where the agents again learn better strategies, but with more inertia (they evaluate and revise their policies once in a while) and in a myopic manner (the decisions are based on current state of population choices and are not so far-sighted). For example, replicator dynamics captures the evolution of population-choice measures in large population games like matching or congestion games with imitation protocols (see e.g., Taylor and Jonker (1978); Hofbauer and Sigmund (1998); Sandholm (2010)).
Several such learning based studies have established the existence of cyclical behavior (Shapley polygons and limit cycles around a Nash equilibrium, see for example, Brown (1951); Shapley (1963); Matsui (1992); Samuelson (1997); Binmore et al. (2001); Boone and Piliouras (2019)). However, this part of the literature predominately focuses on mixed strategies for finite-action games; further, the analysis is derived by assuming the resultant dynamical system to be the solution of an ordinary differential equation or a differential inclusion. Even the recent studies in this direction, that consider some sort of randomness, focus on games with finite actions (e.g., Heinrich et al. (2023) study the dependency on the outcomes of the dynamics based on the order of player-choices). While our study deals with more generic games — also includes the applications with large (compact) action spaces and with potentially discontinuous utility functions (where the best responses may not even exist).
Our work more closely relates to other intricate static notions, where potential outcomes are described using some appropriate mathematical equations/objects, without explicitly describing the dynamics. One such notion namely minimal curb set is introduced by Basu and Weibull (1991) and further related studies are available in Voorneveld et al. (2005); Hurkens (1995). The sink strongly connected components (SCCs) of the better response graph and the evolution of certain Markov chains over such graphs are utilized by Papadimitriou and Piliouras (2018, 2019) to describe the outcomes of the games. We would provide more detailed comparison of these notions with EC after introducing the relevant concepts and notations in Sections 3 and 5; for now, we would like to mention that EC is a far more general notion that encompasses all the above notions and can also be utilized to study the discontinuous games.
In Milionis et al. (2022) the authors discuss (and establish) the insufficiency of Nash Equilibrium in representing the outcome of an arbitrary dynamics related to a finite-action finite-player game in terms of mixed-strategies; and our EC is a step forward in providing alternate outcomes of such games, in fact even for games with large (compact) action sets.
Finally, it has been established that several games mirroring various real-world applications possess equilibrium cycles: recommendation systems in Walunj et al. (2025a), ride-hailing platforms in Walunj et al. (2025b), and supply chains in Wadhwa et al. (2025) are shown to posses equilibrium cycles.
@@ 2 Equilibrium Cycle: Definition and examples
Consider a strategic form game $\big<{\mathcal{N}},({\mathcal{A}}_{i})_{i\in{\mathcal{N}}},$ $({\mathcal{U}}_{i})_{i\in{\mathcal{N}}}\big$>, where for each $i,$ ${\mathcal{A}}_{i}$ is a nonempty set in a metric space.
@@
    [Stability] For any player $i$ and opponent action profile ${\bf a}_{-i}\in{\mathcal{E}}_{-i},$ there exists an action $a_{i}\in{\mathcal{E}}_{i}$ such that U_i(a_i, a_-i) ¿ U_i ( ~a_i, a_-i)  for all actions ~a_i A_i E_i.
@@
    [Unrest] For any action profile $\bf a$$\in{\mathcal{E}}$, there exists a player $i\in{\mathcal{N}},$ and an alternate action $a_{i}^{\prime}\in{\mathcal{E}}_{i}$ such that ${\mathcal{U}}_{i}(a_{i}^{\prime},{\bf a}_{-i})>{\mathcal{U}}_{i}(a_{i},{\bf a}_{-i}),$ and further, U_i(a_i’, a_-i) ¿ U_i ( ~a_i, a_-i)  for all actions ~a_i A_i E_i.
@@
Proposition 1.
For the visibility game discussed in Example 1, the set $0,0.5]^{2}$ is an equilibrium cycle.
@@
If $a_{-i}<0.5,$ then any choice of $a_{i}\in(a_{-i},0.5),$ which implies ${\mathcal{U}}_{i}(a_{i},a_{-i})=1-a_{i},$ satisfies (2). On the other hand, if $a_{-i}=0.5,$ then $a_{i}=0,$ which implies ${\mathcal{U}}_{i}(a_{i},a_{-i})=0.5,$ satisfies (2). This proves that $0,0.5]^{2}$ satisfies the first condition for an EC.
Next, we prove that $0,0.5]^{2}$ satisfies the second condition for an EC. Consider any $(a_{1},a_{2})\in[0,0.5]^{2}.$
  • 1.
@@
This proves that $0,0.5]^{2}$ satisfies the second condition for an EC.
To establish the third condition, suppose, for the purpose of obtaining a contradiction, that $\mathcal{F}$$\subsetneq{\mathcal{E}}$ is a non-empty closed Cartesian product set satisfying the first two conditions for an EC. Define the diagonal ${\mathcal{D}}:=\{(x,x):x\in[0,0.5]\}$. Note that we cannot have $\mathcal{D}$$\subset{\mathcal{F}}$, since this would imply (given that $\mathcal{F}$} is a Cartesian product) that $\mathcal{F}$$\mathcal{E}$}. It follows that ${\mathcal{F}}\cap{\mathcal{D}}\subsetneq{\mathcal{D}}.$ Consider the following cases.
@@ 2. $(ii)$
    $\{x\in[0,b)\colon(x,x)\in{\mathcal{F}}\}\neq\emptyset:$ In this case, define c := sup{x : (x, x) ∈F∩[0,b)^2}. By the construction of $b$ and $c$, it follows that $c<b,$ $(c,c)$ lies in $\mathcal{F}$}, and $(x,x)\notin{\mathcal{F}}$ for all $x\in(c,b).$ Now, for any player $i,$ fix opponent’s action $a_{-i}=c.$ Then for $a_{i}\in(c,b),$ U_i(a_i, c) = 1 - a_i ¿ U_i(~a_i, c) ∀ ~a_i ∈F_i. This contradicts that $\mathcal{F}$} satisfies the first condition for an EC.
In summary, we have shown (via a contradiction-based argument) that there does not exist a closed Cartesian product $\mathcal{F}$$\subsetneq{\mathcal{E}}$ that satisfies the first two conditions of an EC. We conclude that $0,0.5]^{2}$ is an EC for the visibility game in Example 1.
Interestingly, it is known that the visibility game of Example 1 admits a mixed Nash equilibrium supported on $0,1-\nicefrac{{1}}{{e}}]^{2},$ where $e$ denotes Euler’s number; see Lotker et al. (2008). It is instructive to note that the support of the mixed NE differs from the EC, i.e., $0,0.5]^{2}.$
@@
It can be shown that $0,\nicefrac{{(N-1)}}{{N}}]^{N}$ is an EC for this $N$ player extension. Observe that the action set corresponding to each player in this EC grows as $N$ increases.
@@
Consider the following variation of the two-player Bertrand duopoly (see Osborne (2009)) where firms additionally incur a fixed operational cost. Formally, the firms have a fixed marginal cost of production $c>0$, and a fixed operational cost $O_{c}>0$ for participating in the market. Each player/firm $i$ has two options: it can either operate (incurring the cost $O_{c}$) by setting price $p_{i}\geq 0$, or choose not to operate, via the ‘opt-out’ action $p_{i}=n_{o}<0$ (the opt-out action is represented by a negative number for mathematical convenience), incurring no costs. Thus, the action space of each player is modeled as $0,\infty)\cup\{n_{o}\}.$
@@
${\mathcal{E}}:=\left([p_{b},p_{m}^{*}]\cup\{n_{o}\}\right)^{2}$
@@
It is instructive to note that in the classical Bertrand duopoly, $O_{c}=0,$ resulting in the unique Nash equilibrium $(c,c).$ It is the (reasonable) introduction of a positive operational cost (that is not so large that it prohibits firms from operating at all, as ensured by our assumption) that results in the EC characterized in Proposition 2. Finally, we note that the linear demand assumption is not essential to the emergence of the EC; an analogous EC can be established under any smooth decreasing demand function (so long as a positive utility is possible). The proof of Proposition 2 is analogous to that of Proposition 1, and is therefore omitted.
The Bertrand duopoly posits an extreme ‘all or nothing’ bifurcation of market demand between the firms. The following example, which generalizes a pricing game between competing ride-hailing platforms in Walunj et al. (2025b), considers a ‘less extreme’ payoff discontinuity.
@@
The game in Example 3 also admits an EC, as shown in the following lemma; this generalizes Theorem 4 in Walunj et al. (2025b). The proof is once again omitted, given its similarity to the proof of Proposition 1.
Proposition 3.
Consider the game described in Example 3. Let $b:=f^{-1}(g(b))$ and $c:=\operatorname*{arg\,max}_{a_{i}\in{\mathcal{A}}_{i}}g(a_{i}),$ Then ${\mathcal{E}}=[b,c]^{2}$ is an equilibrium cycle of this game.
@@ 3.1 Curb sets in best response games
    To prove that $C$ satisfies the first property of an EC, consider any player $i$ and ${\bf a}_{-i}\in C_{-i}.$ Using the definition of the curb set (see (5)), against ${\bf a}_{-i},$ there exists a best response $a_{i}\in C_{i}$ such that U_i(a_i, a_-i) ¿ U_i(~a_i, a_-i) for all ~a_i ∉C_i (indeed, note that no other outside action $\tilde{a}_{i}\notin C_{i}$ is a best response). Thus, the first property of EC is satisfied.
@@
    Consider any strategy profile ${\bf a}\in C$. Since $C$ is a non-trivial curb set, $\bf a$} is not an NE, and thus there exists at least one player (say player $i$) which stands to obtain a strictly better utility via unilateral deviation. Suppose $a_{i}^{\prime}$ is a best response of this player $i;$ and we have ${\mathcal{U}}_{i}(a_{i}^{\prime},{\bf a}_{-i})>{\mathcal{U}}(a_{i},{\bf a}_{-i})$. Further, by the definition of curb set, $a_{i}^{\prime}\in C_{i}$ and as before, U_i(a_i’, a_-i) ¿ U_i(~a_i, a_-i) for all ~a_i ∉C_i. Thus, $C$ satisfies the second property of an EC.
@@
    We prove that $C$ satisfies the third property of an EC via a contradiction-based argument. Specifically, suppose that there exists ${\mathcal{E}}\subsetneq C$ satisfying the first two properties of an EC. Consider any player $i$. Against any ${\bf a}_{-i}\in{\mathcal{E}}_{-i}$, from the first property of EC, there exists $a_{i}^{\prime}\in{\mathcal{E}}_{i}$ such that U_i(a_i’, a_-i) ¿ U_i ( ~a_i, a_-i) for all ~a_i ∉E_i . This implies that the best responses of player $i$ must be contained in ${\mathcal{E}}_{i},$ which in turn implies that the set $\mathcal{E}$} contains all its best responses. This means that ${\mathcal{E}}\subsetneq C$ is also a curb set, thus contradicting the assumption that $C$ is a minimal curb set.